The speed of yellow light(from a sodium lamp) in a certain liquid is measured to be 1.92*10power8 m/s.What is the index of refraction of this liquid for the light?

Grade:Upto college level

1 Answers

shaik mohammad gouse
10 years ago

SPEED OF YELLOW LIGHT =1.92*10 POWER8 m/s

we know that speed of light = 3*10 power 8

we know that refraction index ( n) =c/v

                                                    =3*10power8/1.92*10power8

                                                   =  1.5625